# Imaging two-body correlations in atomic nuclei via low- and high-energy processes

### v1

@@ Introduction.
Traditionally, the collective behavior of nucleons has been inferred through its fingerprints on low-energy observables. Most notably, subsets of excited states displaying specific energy spectrum and electromagnetic decay probabilities can be associated, through a simplistic model, with a rotor (i.e., rotational excitations) characterized by a given intrinsic deformation or with a harmonic oscillator (i.e., vibrational excitations) characterized by a given phonon frequency. Associated correlations can be quantified via the ground-state expectation value of so called Kumar operators [25, 28] that can, in principle, be reconstructed experimentally by combining electromagnetic transition probabilities to a set of excited states [9, 16, 2].
@@
During the first steps of this endeavor [19, 4, 39, 30, 10], the analyses were performed based on the model picture of the nucleus as a classical rotor, i.e., a rotating deformed rigid body. While this simple interpretation in terms of a model-dependent intrinsic deformation is qualitatively applicable to nuclei whose experimental spectrum display a clear “rotational” band on top of the ground state, it is not adapted to other nuclei, i.e., doubly-magic, semi-magic and so-called transitional nuclei. In spite of the key results such early analyses delivered [4, 39, 30, 17, 21], the situation calls for a more robust interpretation of the data based on a general quantum mechanical description of the colliding nuclei.
@@
Present ab initio nuclear structure calculations are based on the Projected Generator Coordinate Method (PGCM) [15, 14, 20]. This is a variational approach based the many-body ansatz
@@
where the normalization constant $c_{\pm\ell}$ reported in the Supplemental Material (SM) is chosen in a way that ensures $Y_{\ell}^{\pm\ell}(\Omega)$ is normalized on the unit sphere [^1] The spherical harmonic $Y_{\ell}^{\ell}(\Omega_{i})$ with maximum projection $\mu=\ell$, where the quantized z-axis is the axis of the reaction, is nothing but the $\ell$-moment of the eccentricity $(x_{i}+iy_{i})^{\ell}$ in the transverse plane of the reaction. and where the expectation value is taken with respect to the many-body state of interest $|\Psi^{J\Pi}_{k}\rangle$.
@@
$\langle\epsilon^{2}_{\ell}\rangle$ $\equiv\frac{1}{2}\frac{\langle\delta\mathcal{E}^{(1)}_{\ell}\delta\mathcal{E}^{(1)}_{-\ell}\rangle}{\langle R^{(1)}_{\ell}\rangle^{2}}\equiv\frac{1}{2}\frac{\langle\mathcal{E}^{(2)}_{\ell}\rangle-|\langle\mathcal{E}^{(1)}_{\ell}\rangle|^{2}}{\langle R^{(1)}_{\ell}\rangle^{2}}\;,$ (3)
is, to first order, proportional to the statistical variance of the $\ell^{\text{th}}$ Fourier component $\upsilon_{\ell}$ of the azimuthal flow by virtue of the hydrodynamic nature of the QGP. In Eq. (3), the squared eccentricity operator is defined as $\mathcal{E}^{(2)}_{\ell}\equiv\mathcal{E}^{(1)}_{\ell}\mathcal{E}^{(1)}_{-\ell}$ whereas $R^{(1)}_{\ell}\equiv\sum_{i_{1}=1}^{A}(r^{2}_{i_{1}x}+r^{2}_{i_{1}y})^{\ell/2}$ stands for the one-body transverse radius operator. For $J^{\Pi}=0^{+}$ states under present investigation, the mean eccentricity $\langle\mathcal{E}^{(1)}_{\ell}\rangle$ is zero such that second term in the numerator of Eq. (3) drops. As shown in the SM, $\mathcal{E}^{(2)}_{\ell}$ is the sum of a one-body operator $\mathcal{E}^{(2)\,1\text{b}}_{\ell}$ and of a two-body operator $\mathcal{E}^{(2)\,2\text{b}}_{\ell}$ such that the normalized mean-square eccentricity itself decomposes according to
$\langle\epsilon^{2}_{\ell}\rangle$ $\equiv\langle\epsilon^{2}_{\ell}\rangle^{1\text{b}}+\langle\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (4)
@@
$\langle\epsilon^{2}_{2}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{1}{A}+\frac{3}{4\pi}\beta^{2}_{20}\;,\quad\langle\epsilon^{2}_{3}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{16}{3\pi A}+\frac{2048}{245\pi^{3}}\beta^{2}_{30}\;.$ (5)
The two contributions to $\langle\epsilon^{2}_{\ell}\rangle^{\text{RR}}$ behave very differently: while the one-body contribution is equal to $A^{-1}$ and is independent of the deformation parameter, the two-body contribution probing genuine two-body correlations is independent of the mass $A$ but scales with the square of the intrinsic deformation parameter. Based on this model result, a quantity homogeneous to the square of a dimensionless deformation parameter is introduced according to [7]
$\mathcal{B}^{2}_{\ell}(\text{HE})\equiv\frac{4\pi}{3}\langle\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (6)
@@
This operator entertains a close proximity with $\mathcal{E}^{(2)}_{2}$, except that the former involves only protons while the latter involves both protons and neutrons [^2] The other difference is that, contrary to $\mathcal{E}^{(2)}_{2}$, $\mathcal{Q}^{(2)}_{2}$ involves the five components of the spherical tensor $Y^{\mu}_{2}(\Omega_{i})$, $\mu=-2,\ldots,+2$, such that it is a scalar with respect to rotations in three dimension. Interestingly, this would in fact only result in a prefactor difference for $J^{\Pi}=0^{+}$ states if the operators were summing over the same set of nucleons.. Thus, $\mathcal{Q}^{(2)}_{2}$ only probes proton correlations such that its expectation value, presently restricted to $J^{\Pi}=0^{+}$ ground-states, can be expressed in terms of reduced electromagnetic quadrupole transition probabilities according to
@@ Results.
Figure 1: Projected HFB ground-state (a) mean-square quadrupole anisotropy $\langle\epsilon^{2}_{2}\rangle$ and (b) its rescaled two-body contribution $\mathcal{B}^{2}_{2}(\text{HE})$ for all even-even nuclei between carbon and nickel. Proton (neutron) magic numbers are indicated by horizontal (vertical) lines.
Panel (a) of Fig. 1 displays the PHFB ground-state mean-square quadrupole eccentricity $\langle\epsilon^{2}_{2}\rangle$ of all even-even nuclei between carbon and nickel. This quantity is positive but particularly small for nuclei characterized by proton and/or neutron magic numbers. This total mean-square quadrupole eccentricity displays mild fluctuations as a function of proton and neutron numbers and essentially decreases with A. The RR approximation to $\langle\epsilon^{2}_{2}\rangle$ in Eq. (5) suggests that this structureless decreasing behavior is driven by the trivial $1/A$ one-body contribution that dominates in the light nuclei under consideration. This is indeed numerically confirmed for our PHFB calculations in Fig. 2 and in the SM.
Figure 2: Projected HFB ground-state $\langle\epsilon^{2}_{2}\rangle$, $\langle\epsilon^{2}_{2}\rangle^{1\text{b}}$ and $\langle\epsilon^{2}_{2}\rangle^{2\text{b}}$ as a function of AA for all even-even nuclei between carbon and nickel.
The key information about two-body quadrupole correlations lies in the two-body contribution $\langle\epsilon^{2}_{2}\rangle^{2\text{b}}$, also shown in Fig. 2, and whose rescaled value $\mathcal{B}^{2}_{2}(\text{HE})$ [Eq. (6)] is displayed in panel (b) of Fig. 1. One observes more distinctive patterns such as small negative values along the $N$ or $Z=8,20$ semi-magic chains. Along the $N$ and $Z=28$ semi magic chains, $\mathcal{B}^{2}_{2}(\text{HE})$ is negative in doubly magic nuclei but becomes slightly positive in between them, testifying for increased quadrupole correlations. For nuclei with $8<N,Z<20$, especially around 20Ne ($Z=N=10$) large positive values of $\mathcal{B}^{2}_{2}(\text{HE})$ are predicted, signaling strong collective quadrupole correlations. While smaller in absolute than around 20Ne, positive values are predicted systematically between magic numbers.
@@
Eventually, $\mathcal{B}^{2}_{2}(\text{HE})$ always captures both the negative offset due to Pauli’s exclusion principle and the nucleus-dependent positive contribution originating from collective shape fluctuations in addition to the contribution matching the rotor intrinsic deformation, i.e., $\mathcal{B}^{2}_{2}(\text{HE})$ must not be blindly interpreted as the intrinsic deformation [^4] While presently demonstrated using in steps PHFB and PGCM approximations, this statement is valid independently of the many-body method used to solve Schroedinger’s equation.. In fact, additional non-dynamical correlations must be added to the PGCM that can further move $\mathcal{B}^{2}_{2}(\text{HE})$ away from the quantum rotor intrinsic deformation value $\beta^{2}_{20}(\text{dHFB})$. While including such non-dynamical correlations to the PGCM is beyond the scope of the present letter, it can eventually be achieved perturbatively [15, 14, 13].
@@
The present work demonstrates via systematic ab initio nuclear structure calculations of light nuclei that two-particle correlations in the azimuthal distribution of final hadrons emitted in ultra-relativistic ultra-central ion-ion collisions can provide a meaningful imaging of nuclear ground-states. This imaging authorizes a robust interpretation in terms of rotational and vibrational collective correlations. In light nuclei under study, the experimental observable however incorporates a trivial contribution overshadowing the two-body correlations of interest. It thus remains to be seen how combining results from different ion-ion collisions of well-chosen species can isolate genuine two-body correlations.
A future interesting development will consist of targeting specific three-particle correlations to isolate three-nucleon correlations in which additional nuclear structure informations of interest leave their fingerprint.
@@ Acknowledgements.
We thank J.-P. Blaizot and G. Giacalone for stimulating discussions as well as A. Tichai for helping us benchmark our numerical implementation. The calculations were performed using computational resources from CCRT (TOPAZE and IRENE supercomputers). This work has received funding from the European Research Council under the European Union’s Horizon Europe Research and Innovation Programme (Grant Agreement No. 101162059).
Appendix A SUPPLEMENTAL MATERIAL
Appendix B Eccentricity variance
B.1 First quantization formulation
@@
is such that $Y_{\ell}^{\pm\ell}(\Omega)$ is normalized on the unit sphere.
@@
B.2 Decomposition of $\mathcal{E}^{(2)}_{\ell}$
The operator $\mathcal{E}^{(2)}_{\ell}$ is the sum of a one-body operator and a two-body operator, i.e.
@@ B.2 Decomposition of $\mathcal{E}^{(2)}_{\ell}$
B.3 Normalized eccentricity variance
@@
$\langle\epsilon^{2}_{\ell}\rangle$ $\equiv\frac{1}{2}\frac{\langle\delta\mathcal{E}^{(2)}_{\ell}\rangle}{\langle R^{(1)}_{\ell}\rangle^{2}}$ (19)
@@
$R^{(1)}_{2\ell}\equiv\mathcal{E}^{(2)\,1\text{b}}_{\ell}\;.$ (21)
@@
$\langle\epsilon^{2}_{\ell}\rangle_{0^{+}}$ $=\frac{1}{2}\frac{\langle\mathcal{E}^{(2)\,1\text{b}}_{\ell}\rangle}{\langle R^{(1)}_{\ell}\rangle^{2}}+\frac{1}{2}\frac{\langle\mathcal{E}^{(2)\,2\text{b}}_{\ell}\rangle}{\langle R^{(1)}_{\ell}\rangle^{2}}\equiv\langle\epsilon^{2}_{\ell}\rangle^{1\text{b}}+\langle\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (22)
B.4 Second quantization formulation

### v2

@@ Introduction.
Traditionally, the collective behavior of nucleons has been inferred through its fingerprints on low-energy observables. Most notably, subsets of excited states displaying specific energy spectrum and electromagnetic decay probabilities can be associated, through a simplistic model, with a rotor (i.e., rotational excitations) characterized by a given intrinsic deformation or with a harmonic oscillator (i.e., vibrational excitations) characterized by a given phonon frequency. Associated correlations can be quantified via the ground-state expectation value of so called Kumar operators [31, 34, 30] that can, in principle, be reconstructed experimentally by combining electromagnetic transition probabilities to a set of excited states [12, 20, 10, 3, 29, 36].
@@
During the first steps of this endeavor [23, 5, 47, 37, 13], the analyses were performed based on the model picture of the nucleus as a classical rotor, i.e., a rotating deformed rigid body [14]. While this simple interpretation in terms of a model-dependent intrinsic deformation is qualitatively applicable to nuclei whose experimental spectrum display a clear “rotational” band on top of the ground state, it is not adapted to other nuclei, i.e., doubly-magic, semi-magic and so-called transitional nuclei. In spite of the key results such early analyses delivered [5, 47, 37, 21, 25], the situation calls for a more robust interpretation of the data based on a general quantum mechanical description of the colliding nuclei.
@@
Present ab initio nuclear structure calculations are based on the Projected Generator Coordinate Method (PGCM) [19, 18, 24]. This is a variational approach based on the many-body ansatz
@@
where the normalization constant $c_{\pm\ell}$ reported in the Supplemental Material (SM) is chosen in a way that ensures $Y_{\ell}^{\pm\ell}(\Omega)$ is normalized on the unit sphere and where the expectation value is taken with respect to the many-body state of interest $|\Psi^{J\Pi}_{k}\rangle$.
@@
$\langle\delta\epsilon^{2}_{\ell}\rangle$ $\equiv\frac{1}{2}\frac{\langle\delta\mathcal{E}^{(1)}_{\ell}\delta\mathcal{E}^{(1)}_{-\ell}\rangle}{\langle R^{(1)}_{\ell}\rangle^{2}}\equiv\frac{1}{2}\frac{\langle\mathcal{E}^{(2)}_{\ell}\rangle-|\langle\mathcal{E}^{(1)}_{\ell}\rangle|^{2}}{\langle R^{(1)}_{\ell}\rangle^{2}}\;,$ (3)
is, to first order, proportional to the statistical variance of the $\ell^{\text{th}}$ Fourier component $\upsilon_{\ell}$ of the azimuthal flow by virtue of the hydrodynamic nature of the QGP. In Eq. (3), the squared eccentricity operator is defined as $\mathcal{E}^{(2)}_{\ell}\equiv\mathcal{E}^{(1)}_{\ell}\mathcal{E}^{(1)}_{-\ell}$ whereas $R^{(1)}_{\ell}\equiv\sum_{i_{1}=1}^{A}(x^{2}_{i_{1}}+y^{2}_{i_{1}})^{\ell/2}$ stands for the one-body transverse radius operator. For $J^{\Pi}=0^{+}$ states under present investigation, the mean eccentricity $\langle\mathcal{E}^{(1)}_{\ell}\rangle$ is zero such that second term in the numerator of Eq. (3) drops. As shown in the SM, $\mathcal{E}^{(2)}_{\ell}$ is the sum of a one-body operator $\mathcal{E}^{(2)\,1\text{b}}_{\ell}$ and of a two-body operator $\mathcal{E}^{(2)\,2\text{b}}_{\ell}$ such that the normalized mean-square eccentricity itself decomposes according to
$\langle\delta\epsilon^{2}_{\ell}\rangle$ $\equiv\langle\delta\epsilon^{2}_{\ell}\rangle^{1\text{b}}+\langle\delta\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (4)
@@
$\langle\delta\epsilon^{2}_{2}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{1}{A}+\frac{3}{4\pi}\beta^{2}_{20}\;,\quad\langle\delta\epsilon^{2}_{3}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{16}{3\pi A}+\frac{2048}{245\pi^{3}}\beta^{2}_{30}\;.$ (5)
The two contributions to $\langle\delta\epsilon^{2}_{\ell}\rangle^{\text{RR}}$ behave very differently: while the one-body contribution is equal to $A^{-1}$ and is independent of the deformation parameter, the two-body contribution probing genuine two-body correlations is independent of the mass $A$ but scales with the square of the intrinsic deformation parameter. Based on this model result, a quantity homogeneous to the square of a dimensionless deformation parameter is introduced according to [8]
$\mathcal{B}^{2}_{\ell}(\text{HE})\equiv\frac{4\pi}{3}\langle\delta\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (6)
@@
This operator entertains a close proximity with $\mathcal{E}^{(2)}_{2}$, except that the former involves only protons while the latter involves both protons and neutrons. [^1] The other difference is that, contrary to $\mathcal{E}^{(2)}_{2}$, $\mathcal{Q}^{(2)}_{2}$ involves the five components of the spherical tensor $Y^{\mu}_{2}(\Omega_{i})$, $\mu=-2,\ldots,+2$, such that it is a scalar with respect to rotations in three dimension. Interestingly, this would in fact only result in a prefactor difference for $J^{\Pi}=0^{+}$ states if the operators were summing over the same set of nucleons. Thus, $\mathcal{Q}^{(2)}_{2}$ only probes proton correlations such that its expectation value, presently restricted to $J^{\Pi}=0^{+}$ ground-states, can be expressed in terms of reduced electromagnetic quadrupole transition probabilities according to
@@ Results.
Figure 1: Projected HFB ground-state (a) mean-square quadrupole anisotropy $\langle\delta\epsilon^{2}_{2}\rangle$ and (b) its rescaled two-body contribution $\mathcal{B}^{2}_{2}(\text{HE})$ for even-even nuclei between carbon and nickel. Proton (neutron) magic numbers are indicated by horizontal (vertical) lines.
Panel (a) of Fig. 1 displays the PHFB ground-state normalized mean-square quadrupole eccentricity $\langle\delta\epsilon^{2}_{2}\rangle$ of even-even nuclei between carbon and nickel. This quantity is positive but particularly small for nuclei characterized by proton and/or neutron magic numbers. This total normalized mean-square quadrupole eccentricity displays mild fluctuations as a function of proton and neutron numbers and essentially decreases with A. The RR approximation to $\langle\delta\epsilon^{2}_{2}\rangle$ in Eq. (5) suggests that this structureless decreasing behavior is driven by the trivial $1/A$ one-body contribution that dominates in the light nuclei under consideration. This is indeed numerically confirmed for our PHFB calculations in Fig. 2 and in the SM.
Figure 2: Projected HFB ground-state $\langle\delta\epsilon^{2}_{2}\rangle$, $\langle\delta\epsilon^{2}_{2}\rangle^{1\text{b}}$ and $\langle\delta\epsilon^{2}_{2}\rangle^{2\text{b}}$ as a function of AA for even-even nuclei between carbon and nickel.
The key information about two-body quadrupole correlations lies in the two-body contribution $\langle\delta\epsilon^{2}_{2}\rangle^{2\text{b}}$, also shown in Fig. 2, and whose rescaled value $\mathcal{B}^{2}_{2}(\text{HE})$ [Eq. (6)] is displayed in panel (b) of Fig. 1. One observes more distinctive patterns such as small negative values along the $N$ or $Z=8,20$ semi-magic chains. Along the $N$ and $Z=28$ semi magic chains, $\mathcal{B}^{2}_{2}(\text{HE})$ is negative in doubly magic nuclei but becomes slightly positive in between them, testifying for increased quadrupole correlations. For nuclei with $8<N,Z<20$, especially around 20Ne ($Z=N=10$) large positive values of $\mathcal{B}^{2}_{2}(\text{HE})$ are predicted, signaling strong collective quadrupole correlations. While smaller in absolute than around 20Ne, positive values are predicted systematically between magic numbers.
@@
Eventually, $\mathcal{B}^{2}_{2}(\text{HE})$ always captures both the negative offset due to Pauli’s exclusion principle and the nucleus-dependent positive contribution originating from collective shape fluctuations in addition to the contribution matching the rotor intrinsic deformation, i.e., $\mathcal{B}^{2}_{2}(\text{HE})$ must not be blindly interpreted as the intrinsic deformation. [^3] This statement is valid independently of the many-body method used to solve Schrödinger’s equation.. In fact, missing dynamical correlations must be added to the PGCM that can further move $\mathcal{B}^{2}_{2}(\text{HE})$ away from the quantum rotor intrinsic deformation value $\beta^{2}_{20}(\text{dHFB})$. While including such dynamical correlations to the PGCM is beyond the scope of the present letter, it can eventually be achieved perturbatively [19, 18, 17]
@@
Finally, one notices that $\mathcal{B}^{2}_{2}(\text{HE})$, which relates to the two-body part of $\langle\delta\epsilon^{2}_{2}\rangle$ through Eq. (6), is much larger in 20Ne than in 16O. Given that the one-body contribution to $\langle\delta\epsilon^{2}_{2}\rangle$ is very similar in both nuclei, the excess of elliptic flow in 20Ne + 20Ne collisions compared to 16O + 16O, initially predicted in Refs. [21, 25] and observed experimentally at the LHC [27, 2, 1], directly reflects the different nature of their two-body quadrupolar correlations.
@@
The present work demonstrates via systematic ab initio nuclear structure calculations of light nuclei that two-particle correlations in the azimuthal distribution of final hadrons emitted in ultra-relativistic ultra-central ion-ion collisions can provide a meaningful imaging of nuclear ground-states. This imaging authorizes a robust interpretation in terms of rotational and vibrational collective correlations. In light nuclei under study, the experimental observable however incorporates a trivial contribution overshadowing the two-body correlations of interest. It is thus necessary to combine results from different ion-ion collisions of well-chosen species to isolate, or at least magnify, genuine two-body correlations. This is what has been already achieved recently by taking the ratio of the 20Ne + 20Ne over 16O + 16O elliptic (triangular) flows to magnify quadrupolar (octupolar) correlations at play in both nuclei.
A future interesting development will consist of targeting specific three-particle correlations to isolate three-nucleon correlations in which additional nuclear structure information of interest leave their fingerprint.
@@ Acknowledgements.
We thank J.-P. Blaizot and G. Giacalone for stimulating discussions as well as V. Somà for proofreading the manuscript. We also thank A. Tichai for helping us benchmark our numerical implementation. The calculations were performed using computational resources from CCRT (TOPAZE and IRENE supercomputers). This work has received funding from the European Research Council under the European Union’s Horizon Europe Research and Innovation Programme (Grant Agreement No. 101162059).
I Supplemental Material
II Eccentricity variance
II.1 First quantization formulation
In the following, the position of a nucleon will be labeled by the spherical coordinates, $r$ and $\Omega$, as well as the cartesian coordinates, $x$ and $y$.
@@ II.1 First quantization formulation
is such that $Y_{\ell}^{\pm\ell}(\Omega)$ is normalized on the unit sphere. Introducing for later use the one-body spherical tensor operators $E^{\pm\,(1)}_{\ell}\equiv\{E^{\pm\,(1)}_{\ell\mu},\mu=-\ell,-\ell+1,...,\ell-1,\ell\}$ whose components are defined as
$E^{\pm\,(1)}_{\ell\mu}$ $=\sum_{i_{1}=1}^{A}\frac{r_{i_{1}}^{\ell}}{c_{\pm\ell}}\,Y_{\ell}^{\mu}(\Omega_{i_{1}})\;,$ (14)
the eccentricity operators are nothing but
$\mathcal{E}^{(1)}_{\pm\ell}\equiv E^{\pm\,(1)}_{\ell\pm\ell}\;.$ (15)
@@
II.2 Decomposition of $\mathcal{E}^{(2)}_{\ell}$
The operator $\mathcal{E}^{(2)}_{\ell}$ is the sum of a one-body operator and a two-body operator, i.e.,
@@ II.2 Decomposition of $\mathcal{E}^{(2)}_{\ell}$
II.3 Normalized eccentricity variance
@@
$\langle\delta\epsilon^{2}_{\ell}\rangle$ $\equiv\frac{1}{2}\frac{\langle\delta\mathcal{E}^{(2)}_{\ell}\rangle}{\langle R^{(1)}_{\ell}\rangle^{2}}$ (21)
@@
$\mathcal{E}^{(2)\,1\text{b}}_{\ell}\equiv R^{(1)}_{2\ell}\;$. (23)
@@
$\langle\delta\epsilon^{2}_{\ell}\rangle_{0^{+}}$ $=\frac{1}{2}\frac{\langle\mathcal{E}^{(2)\,1\text{b}}_{\ell}\rangle_{0^{+}}}{\langle R^{(1)}_{\ell}\rangle^{2}_{0^{+}}}+\frac{1}{2}\frac{\langle\mathcal{E}^{(2)\,2\text{b}}_{\ell}\rangle_{0^{+}}}{\langle R^{(1)}_{\ell}\rangle^{2}_{0^{+}}}$
$\equiv\langle\delta\epsilon^{2}_{\ell}\rangle^{1\text{b}}_{0^{+}}+\langle\delta\epsilon^{2}_{\ell}\rangle^{2\text{b}}_{0^{+}}\;.$ (24)
II.4 Second quantization I